28) A sample of sulfur contains 1.81 x 1024 atoms of sulfur How many moles of sulfur 28) does this represent? A) 3.01 mol B) 1.81 mol C) 1.77 mol D) 1.09 mol E) 0.333 mol 29) When the equation below is balanced with the lowest possible whole number coefficients, what is the sum of the coefficients? 29) Ca(s) + HCl(r) → CaCl2(s) + H26) A) 2 B) 5 C) 4 D) 8 E) 10 o) Aspirin is the common name for acetyl salicylic acid, CoHg04. One tablet contains 0.325 g of aspirin. How many moles of aspirin is this? 30) A) 1.80 x 10-3 mol B) 0.554 mol C) 1.96 × 1023 mol D) 1.80 x 10-6 mol E) 554 mol 31) Which statement concerning the three states of matter is FALSE? Particles of a solid have low kinetic energy, and are held close together by strong attractive forces A) B) Particles of a gas are well separated, have no regular arran C) Particles of a liquid have a ngid, ordered arra D) Phase changes E) Gases have low 31) gement, and rm ove at high speeds the liquid particles in the identity or composition of the substance occurs ngement due to the high speeds of es between the three states of matter are physical changes; no change densities and high compressibility due to the large amount of space between gas particles 32) A) expansion of gas particles in an open container B the resistance of gas particles to flow C) collisions of gas particles with each other D) collisions of gas particles with the walls of the conta is a result of which of the following? )Gas pressure i E) compressibility of gas particles
Answer #1

28)

use:

number of mol = number of atoms / Avogadro’s number

number of mol = 1.81*10^24 / 6.022*10^23 mol

number of mol = 3.01 mol

Answer: 3.01 mol

29)

The given reaction is:

Ca(s) + HCl (g) —> CaCl2 (s) + H2 (g)

Balance Cl:

Ca(s) + 2 HCl (g) —> CaCl2 (s) + H2 (g)

The equation is now balanced

Sum of coefficient = 1 + 2 + 1 + 1 = 5

Answer: B

30)

Molar mass of C9H8O4,

MM = 9*MM(C) + 8*MM(H) + 4*MM(O)

= 9*12.01 + 8*1.008 + 4*16.0

= 180.154 g/mol

mass(C9H8O4)= 0.325 g

use:

number of mol of C9H8O4,

n = mass of C9H8O4/molar mass of C9H8O4

=(0.325 g)/(1.802*10^2 g/mol)

= 1.80*10^-3 mol

Answer: A

Only 1 question at a time please